OpenStudy (anonymous):

An arrow is shot into the air from the top of a castle wall, which has a height of 98 ft, into the fields surrounding the castle. It is shot with the initial velocity of 54 ft/s, and it reaches its maximum height when it is 107 ft from the base of the castle wall. How many seconds will it take to land? How far will it land from the base of the castle wall? Give your answer correct to the nearest tenth of a unit [HINT: Use the symmetry of the parabola and acceleration due to gravity g = -32 ft/s^2

OpenStudy (anonymous):

Hi there. Do you just want an answer or do you want to learn the material?

OpenStudy (anonymous):

i would like to know the material please

OpenStudy (anonymous):

to do this use the equations of constant acceleration for vertical component of the motion and v = dist/time for horizontal component

OpenStudy (anonymous):

im in calculus

OpenStudy (anonymous):

i need to do it the calculus way

OpenStudy (anonymous):

OK, wonderful. Well the first thing to remember is that the arrow is fired at some angle. Let's call it a. And at the instant it leaves the archer's bow, the total velocity can be divided up into two parts: it's vertical velocity, and its horizontal velocity. The way it's divided up between the two depends on the angle. At high angles, almost all the velocity is vertical. At low angles, it's almost all horizontal. Here's the thing: the horizontal velocity will remain the same for the whole flight, but the vertical velocity will become less and less (under the force of gravity), eventually become negative, then get faster and faster until the arrow strikes the ground. Does that make sense so far?

OpenStudy (anonymous):

uh huh

OpenStudy (anonymous):

OK, so first, using trig, what's the equation for the vertical component of the arrow's velocity, in terms of its initial velocity (54) and its firing angle (a)?

OpenStudy (anonymous):

is this the physics way?

OpenStudy (anonymous):

no it's the calculus way

OpenStudy (anonymous):

it's sort of both, physics and calculus go hand in hand, but i'm not assuming any prior knowledge of the linear motion equations as a physics student would

OpenStudy (anonymous):

so i would have a(t) = -32 v(t) = -32t+54

OpenStudy (anonymous):

OK, so first, using trig, what's the equation for the vertical component of the arrow's velocity, in terms of its initial velocity (54) and its firing angle (a)?

OpenStudy (anonymous):

im not sure..

OpenStudy (anonymous):

s(t)=-16t^2+54t+C

OpenStudy (anonymous):

i have 4.7 sec takes to land is that right?

OpenStudy (anonymous):

Hi I finally figured this stupid thing out too, haha

OpenStudy (anonymous):

I set up an equation that related x (distance from wall) to y (height above ground), with three unknown constants (a, b, and c): h(x) = ax^2 + bx + c, I simultaneously solved: 1) h[0] = 98 (the arrow starts out at 98 feet above) 2) h'[107] = 0 (the arrow's vertical velocity when it's 107 feet away is 0) 3) h''[0] = -32 (the arrow's vertical acceleration is -32feet/second squared) These three equations, when you differentiate and substitute, come out to: 1) c = 98 2) 214a + b = 0 3) 2a = -32 Simultaneously solve and you get a = -16, b = 3424, c = 98.

OpenStudy (anonymous):

Now if you want to know where it lands, you solve for 0: -16x^2 + 3424x + 98 = 0 x = 214.029 So it lands 214.029 feet from the castle wall.

OpenStudy (anonymous):

i came up with the formula -16x^2+54X+98 how do i find the distance from that?

OpenStudy (anonymous):

OK I just figured out why this problem was giving me so much trouble. It's impossible. The problem is malformed. If you shoot the arrow straight up, it will reach its maximum height in 54/32 seconds, or about 1.69 seconds. That's the time it takes gravity to stop an arrow shot at 54 m/s. At any angle other than straight up, the arrow will reach maximum height even sooner; that's the maximum time between firing and max height. However, even if the arrow were moving directly horizontally for 1.69 seconds, it would only go 91.1 feet horizontally -- short of the 107 feet from the castle base this problem claims. It's impossible to fire an arrow at 54 meters/second at any angle under the influence of 32m/s/s gravity and have it be 107 feet from the castle base at its maximum height.

OpenStudy (anonymous):

ok thanks a lot! so i can't figure this problem out even if i try to?

OpenStudy (anonymous):

no, the problem is malformed